OpenStudy (shaima):

Will Fan and medal Hera spent $16 on dinner. Then she spent $6 on dessert and some more on a taxi ride. Including the taxi fare, she spent a total of $31. How much money did Hera spend on the taxi ride?

OpenStudy (shaima):

(a) What is the unknown information? (b) Let x represent the unknown information. Write an addition equation that models the problem. (c) Solve the equation. (d) Answer the question.

OpenStudy (brainlesslonelygirl):

the point you've mentioned should help you solve the question. try it this way, you'll get the total cost if you add all the money she spent right?

OpenStudy (brainlesslonelygirl):

so if you take x as the money she spent on taxi. do you think you can make an equation?

OpenStudy (shaima):

i think

OpenStudy (shaima):

16+6=22

OpenStudy (shaima):

so 9 dollars on the taxi ride?

OpenStudy (brainlesslonelygirl):

16+6= 22 now that is the money for dinner and dessert. what about the taxi money?

OpenStudy (shaima):

22+9 is 31

OpenStudy (brainlesslonelygirl):

sorry i read a number wrong lol yes the money on taxi is 9 good job but its important you know how to make an equation in x to solve for the answer

OpenStudy (shaima):

so 16+6=(22)=9=31

OpenStudy (shaima):

(22)+9*

OpenStudy (brainlesslonelygirl):

basically. 16+6+x=31 then you solve for x

OpenStudy (shaima):

22?

OpenStudy (brainlesslonelygirl):

yes when you plus 16+6 thats 22 if you want you can add those two first then add x and equate to 31 22+x=31

OpenStudy (shaima):

(a) What is the unknown information? (b) Let x represent the unknown information. Write an addition equation that models the problem. (c) Solve the equation. (d) Answer the question. but we still need to answer this

OpenStudy (brainlesslonelygirl):

a. unknown information is the taxi money b. equation is 16+6+x=31 c.solving the equation x=31-16+6 so x= 9 d. hera spent $9 on taxi

OpenStudy (shaima):

ty

OpenStudy (brainlesslonelygirl):

anytime(=